OpenStudy (anonymous):

Which property states that the order of numbers can be changed in an addition or multiplication problem? A. Associative B. Commutative C. Transitive D. Reflexive

OpenStudy (anonymous):

Commutative Property.. 1. In Addition: \[(a + b) = (b +a)\] 2. In Multiplication: \[(a \times b) = ( b \times a)\]

OpenStudy (anonymous):

so what letter is it?